sketch in the normal table and determine the probability of the following z scores

a. p(z<+0.85)

b. p(z > - 1.23)

c. p(-1.50 <z< + 1.50)

d. p(0.00 <z< + 2.27)


1
Expert's answer
2021-05-11T12:25:08-0400

The probability  is equal to the blue area under the curve.

Using the standard normal table we find this area.



a)


P(Z<0.85) = 0.8023

b)


P(Z>-1.23) = 0.8907

c)


To find the probability of P(-1.5<Z<1.5), we use the following formula:

P(-1.5<Z<1.5) = P(Z<1.5) - P(Z< -1.5)

P(Z<1.5) can be found from a standard normal table, an it is equal to 0.9332

P(Z< -1.5) = 1 - P(Z<1.5) = 1 - 0.9332 = 0.0668

So,

P(-1.5<Z<1.5) = 0.9332 - 0.0668 = 0.8664

d)


P(0<Z<1.5) = 0.4884